40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 03:08:57 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec DRFPMI_locked.pdf
    Reply  Thu Oct 15 04:14:14 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Thu Oct 15 10:59:12 2015, Koji, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Sat Oct 17 18:55:07 2015, rana, Update, LSC, DRFPMI Locked for 20 sec 
       Reply  Mon Oct 19 15:23:22 2015, ericq, Update, LSC, Longer DRFPMI lock second_DRFPMI_15Oct2015.png
          Reply  Tue Oct 20 11:24:29 2015, ericq, HowTo, LSC, How to DRFPMI 
Message ID: 11696     Entry time: Sat Oct 17 18:55:07 2015     In reply to: 11691     Reply to this: 11698
Author: rana 
Type: Update 
Category: LSC 
Subject: DRFPMI Locked for 20 sec 

smiley in addition to Koji's words I feel like we should also thank those who made small but positive contributions. Its hard not to notice that this locking only happened after the new StripTool PEM colors were implemented...

From the times series plot I guess that the fuzz of the in-loop DARM is 1 pm RMS (based on memory). This means that the ALS was holding the DARM at 10 pm from the RF resonances.

There is no significant shift in the DRMI error signals, so new weird CARM effect. Would be interesting to see what the 1f signals do in the last 60 seconds before RF lock.

For documentation, perhaps Gautam can post the loop gain measurements of the 5 loops on top of the Bode plots of the loop models.

ELOG V3.1.3-